t Banked Frictionless Curve, and Flat Curve with FPLEASE ANSWER PART B. THANKS!
Answer #1


if theta = 0 degrees

then v^2 = mu*rg

coefficient of static friction is mu_min = v^2/(r*g)

v = 45 km/hr = 45*5/18 = 12.5 m/sec

mu_min = 12.5^2/(43.8*9.8) = 0.364